1. Design an automaton to accept the language ((a²b3)* :k:0})*
Answer #1

1. Design an automaton to accept the language ((a²b3)* :k:20}).

Please refer to below image for DFA:

Accepting state a Start a q1 92 q3 b b a b a 96 94 a,b a 95 b b DFA

1. Design an automaton to accept the language ((a²b3)* :k:20}).

The language should accepts strings of form

L = {empty string, aabbb, aabbbaabbbb, aabbbaabbbaabbb, ... }

  • q1 : Start state, also accepting state, strings of form  ((a^{2}b^{3}))^{k}: k\geq 0
  • q2:  strings of form  ((a^{2}b^{3}))^{k}a: k\geq 0
  • q3:  strings of form  ((a^{2}b^{3}))^{k}aa: k\geq 0
  • q4:  strings of form  ((a^{2}b^{3}))^{k}aab: k\geq 0​​​​​​​​​​​​​​
  • q5:  strings of form  ((a^{2}b^{3}))^{k}aabb: k\geq 0​​​​​​​​​​​​​​
  • q6: random strings

-------------------------------------

I hope this helps you,

Please rate this answer if it helped you,

Thanks for the opportunity